OpenStudy (anonymous):

3x times (square root 54x) + 2 times (square root 6x^3) - square root 216x^3

OpenStudy (anonymous):

3x√(54x) + 2√(6x^3) - √(216x^3)

OpenStudy (anonymous):

OK the way you do this is you factor the terms inside the square roots as to get perfect squares. I'll do the first one for you: 3x√(54x) = 3x√(9*6x) = 3x√9√(6x) = 3x*3√6x=9x√6x Understand how I got that?

OpenStudy (anonymous):

how did you simplify the 216x^3? that was the only one i had a problem with

OpenStudy (anonymous):

Can you factor 216?

OpenStudy (anonymous):

It's not an easy one. But try making a prime factorization tree if you're stuck.

OpenStudy (anonymous):

wait i got it! thank you!

OpenStudy (whpalmer4):

@BluFoot, if you write sqrt{expr} instead of sqrt(expr) the radical sign extends over the whole expression \[\sqrt{216x^3}= \sqrt{2*2*2*3*3*3*x*x*x}\]

OpenStudy (anonymous):

Hahaha I didn't do much :P good job! your final answer should be 5*x^(3/2)*√6

OpenStudy (anonymous):

@whpalmer4 ooh let me try that 3xsqrt{54x} + 2sqrt{6x^3} - sqrt{216x^3}

OpenStudy (anonymous):

well that didn't go too well

OpenStudy (whpalmer4):

Sorry, that's within the backslash openbracket backslash closebracket \ [ \sqrt{216x^3} \ ] remove the spaces and you get \[\sqrt{216x^3}\]

OpenStudy (anonymous):

Ah. Yeah i'm usually too lazy to do that, unless it's a complex equation. The equation editor is very clumsy in my opinion, they should simplify it. All those square brackets and backward slashes are totally unnecessary.

OpenStudy (whpalmer4):

I don't use the equation editor most of the time, just type it in directly and swear a lot :-)

OpenStudy (anonymous):

I just have fancy keys mapped to my keyboard when I hold shift and alt. So I can do stuff like this with a couple key presses: ∆≈∫µ∑√Ω≤≥π

OpenStudy (whpalmer4):

Ah, that explains it. You're actually inserting a radical sign glyph, not telling it to draw a radical sign around the following expression. Quick and easy, unless you need something like \[\sqrt[3]{\frac{x^2-9}{\frac{x+1}{x-1}}}\]

OpenStudy (whpalmer4):

Well, I'm going to stop hijacking this thread now :-)